OpenStudy (wach):

How does 3^x + 3^x + 3^x = 3^x+1 ?

OpenStudy (asnaseer):

it is making use of the rule:\[a\times a^n=a^{n+1}\]

OpenStudy (asnaseer):

can you see how?

OpenStudy (wach):

Not really. Could you put the equation into that rule? I think I'l be able to understand, then.

OpenStudy (asnaseer):

ok, first of all how can you simplify the left hand side of your expression?

OpenStudy (asnaseer):

e.g. a + a + a = ?

OpenStudy (wach):

3(3^x), right?

OpenStudy (asnaseer):

correct

OpenStudy (asnaseer):

now apply the rule

OpenStudy (waleed_imtiaz):

Lets see this question like this.... :) let 3^x=y..... the above equation is....... 3^x + 3^x + 3^x = 3^x *3 y+y+y=y(3) 3y=3y...... It is an identity...... :DA

OpenStudy (wach):

a(a^n) = a^n+1 3(3^x) = 3^x+1 Aa. If I may ask, how does the rule work/make sense?

OpenStudy (asnaseer):

e.g.:\[a\times a^3=a\times(a\times a\times a)=a\times a\times a\times a=a^4\]

OpenStudy (asnaseer):

and 4 = 3 + 1

OpenStudy (asnaseer):

the general rule here is:\[a^m\times a^n=a^{m+n}\]

OpenStudy (wach):

Ohhhh. That makes complete sense when worked out. Thank you so much for helping me with this concept, asnaseer!

OpenStudy (asnaseer):

yw :)